Mathematics 242 – Principles of Analysis Definitions Quiz 4 – March 1, 2013

Your Name:

A) (5) Define: The real number a is lub(A) for a subset A of the real numbers.

B) (5) Define: The sequence {x_{n}} converges to a ∈ R.